# GRU: Mitigating the Trade-off between Unlearning and Retention for Large Language Models

### v1

@@ 3.3 Theoretical Analysis
Assume the unlearning objective $\mathcal{L}$ is differentiable, $L$-smooth, and lower bounded. Then, the GRU update rule with the learning rate $\texttt{lr}<{2}/{L}$ will converge to either a) a degenerate configuration where $\cos\bigl{(}\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\bigr{)}=-1$ at a specific step $t$, or b) the locally optimal solution $\boldsymbol{\theta}^{*}$ that minimizes $\mathcal{L}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})$.
Remark 3.2.
Overall, Theorem 3.1 demonstrates that, from a convergence perspective, the GRU does not compromise the original goal of unlearning. This is contingent upon avoiding those cases where $\cos\bigl{(}\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\bigr{)}=-1$. Moreover, since stochastic optimization is employed for the LLM unlearning process, we can simply overcome this issue by randomly selecting a new data batch from the unlearning dataset, which will then allow the unlearning process to continue to be carried out. Please refer to Appendix A.2 for the detailed proof.
@@
Assume that the retain loss $\mathcal{R}$ is differentiable and $L$-smooth, and the lr-curvature $\mathfrak{H}_{\texttt{lr}}({\mathcal{R}};\boldsymbol{g})$ for $\mathcal{R}$ (cf., Definition A.1) satisfies $\mathfrak{H}_{\texttt{lr}}({\mathcal{R}};\boldsymbol{g})\geq\ell\|\boldsymbol{g}\|^{2}$ for any gradients $\boldsymbol{g}$ and some constant $\ell\leq L$. Let $\boldsymbol{\theta}^{(t+1)}_{\mathrm{gru}}$ and $\boldsymbol{\theta}^{(t+1)}_{\mathrm{u}}$ be the parameters after applying one step of gradient updates for the original $\boldsymbol{\theta}^{(t)}$ with and without GRU, respectively. Then, we can ensure $\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta}^{(t+1)}_{\mathrm{gru}})\;\leq\;\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta}^{(t+1)}_{\mathrm{u}})$ if a) $\ell\geq\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}+\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}L$ and b) $\ell\geq\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}+\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}L+2/\texttt{lr}$.
Remark 3.4.
In heuristics, $\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}+\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}$ measures the degree of conflict between $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ and $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, of which the larger values generally indicate a greater destructive effect on the retain performance. Therefore, condition a) implies that when the degree of conflict is more severe, our requirements for models, characterized by $\ell/L$, will also be more stringent; condition b) states that to ensure the overall efficacy of our proposed method in terms of retention, the learning rate adopted in GRU should not be excessively small. Please refer to Appendix A.3 for the detailed proof.
@@ 4 Go Beyond GRU
Most unlearning methods, including our GRU, rely on retain data to preserve the overall performance. However, retain data can often be biased, motivating us to investigate a challenging scenario where we need to rely exclusively on the unlearn data $\mathcal{D}_{\mathrm{u}}$, without further access to the retain data $\mathcal{D}_{\mathrm{r}}$. To adapt for this setup, we make several adjustments for GRU and propose task vector rectified unlearning (TRU).
@@
The key insight behind TRU is that unlearning typically involves a series of data points rather than a single instance. Thus, for each individual data point $\boldsymbol{s}_{\mathrm{u}}\in\mathcal{D}_{\mathrm{u}}$ targeted for unlearning, the remaining data points within $\mathcal{D}_{\mathrm{u}}$, i.e., $\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\}$, can offer information for retention if used properly. Here, we incorporate the so-called task vectors [27], which are as the essential component of our algorithmic design, into our process.
@@
Task Vector. A task vector typically represents the necessary adjustments for model parameters to incorporate new knowledge. For example, when we want the model to learn from a specific data point $\boldsymbol{s}$, we initiate by fine-tuning the current model parameterized, denoted by $\boldsymbol{\theta}_{\mathrm{org}}$. It can be achieved through $T$ iterations of gradient updates, following $\boldsymbol{\theta}^{(t+1)}=\boldsymbol{\theta}^{(t)}+\texttt{lr}\cdot\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{s};\boldsymbol{\theta}^{(t)})$ with $\boldsymbol{\theta}^{(0)}=\boldsymbol{\theta}_{\mathrm{org}}$ and $\boldsymbol{\theta}_{\boldsymbol{s}}=\boldsymbol{\theta}^{(T)}$. Obviously, $\boldsymbol{T}_{\boldsymbol{s}}$ allows for the augmentation of the original model with the knowledge acquired from $\boldsymbol{s}$ by applying $\boldsymbol{\theta}_{\mathrm{org}}+\boldsymbol{T}_{\boldsymbol{s}}$. Conversely, to unlearn a data point $\boldsymbol{s}_{\mathrm{u}}$, we can reverse this process by subtracting the task vector via $\boldsymbol{\theta}_{\mathrm{org}}-\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$ [28].
@@
Rectified Task Vector. However, the task vector $\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$ still faces the trade-off between unlearning and retention. Hence, a similar constrained updating rule, as outlined in Eq. (7), can be adopted and further adjusted as:
@@
$&\min_{\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}}~{}\|\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}-{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}\|^{2}\\ ~{}\text{s.t.}~{}\langle-\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}&,\nabla_{\boldsymbol{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\};\boldsymbol{\theta}_{\mathrm{org}})\rangle\geq 0.\end{aligned}$ (11)
@@
It mandates that the task vectors be rectified to have no negative impact on other data points. For this purpose, we utilize other unlearning data, i.e., $\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\}$, as reference. Similar to Eq. (8), we have its closed-form solution as
@@
$\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}={\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}+\frac{[\langle{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}},\,\nabla_{\boldsymbol{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\};\boldsymbol{\theta}_{\mathrm{org}})\rangle]_{-}}{\|\nabla_{\boldsymbol{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\};\boldsymbol{\theta}_{\mathrm{org}})\|^{2}}\,{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}.$ (12)
@@
For each data point $\boldsymbol{s}_{\mathrm{u}}$ within $\mathcal{D}_{\mathrm{u}}$, we can calculate the associated task vector $\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}$. Moreover, to ensure standardised influence across data points, these task vectors are further normalized so that their magnitudes are equal to 1. The resulting normalized vectors are denoted as $\bar{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}$, facilitating equitable integration across data points, as shown below.
To unlearn the entire set $\mathcal{D}_{\mathrm{u}}$, we subtract the average of all normalized task vectors from the original model. This is implemented using the formula of
$\boldsymbol{\theta}_{\mathrm{org}}-\frac{\texttt{stg}}{n}\sum_{\boldsymbol{s}_{\mathrm{u}}\in\mathcal{D}_{\mathrm{u}}}\bar{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}},$ (13)
@@
where $n$ represents the number of data points within $\mathcal{D}_{\mathrm{u}}$ and stg indicates the strength of task vector-based unlearning. It ensures a reliable removal of targeted data, meanwhile mitigating the compromise to the overall performance. Moreover, when the number of data points $n$ is substantial, an effective strategy involves randomly dividing the entire unlearning set into several smaller batches. Each batch then serves as a substitute for $\boldsymbol{s}_{\mathrm{u}}$ in Eq. (11), reducing the demands associated with calculating the task vectors.
@@ 5 Experiments
Benchmarks. Our evaluations adopt three representative benchmarks: TOFU [14], WMDP [18], and MUSE [19]. TOFU comprises 200 synthetic author profiles, a total of 4,000 question-answer pairs. It covers different unlearning setups with varying proportions of data targeted to be unlearned, including 1%, 5%, or 10% of the profiles as unlearning sets. WMDP collects a set of sensitive knowledge encountered in practice, further categorized into three areas as biosecurity, cybersecurity, and chemical security. MUSE constructs their unlearning sets using news articles and books, primarily focusing on addressing copyright issues within existing LLMs.
@@
Baselines and Backbones. For the baseline methods, we focus on a set of objective-based approaches, including GA, GD, NPO, and weighted gradient ascent (WGA) [24]. All of these methods have demonstrated their practical significance and are thus adopted in our experiments. Moreover, for the backbone models, we adhere to the default suggestions for each benchmark. We use further fine-tuned LLaMA2-7B-chat [29] and Phi-1.5 [30] for TOFU; Zephyr-7B-beta [31] for WMDP; ICLM-7B [32] for MUSE.
@@
Hyper-parameters Configurations. In our experiments, we use AdamW optimizer [33] with a batch size of $32$ and learning rates $2\times 10^{-5}$ for Phi-1.5 and $1\times 10^{-5}$ for LLaMA2-7B-chat in TOFU; $1\times 10^{-5}$ in MUSE; $4\times 10^{-6}$ in WMDP. Furthermore, the training steps are set to $5$ epochs for TOFU, $1$ epoch for MUSE, and $20$ steps for WMDP. For hyperparameters in GRU, we employ grid search on validation data to identify their optimal values. The candidate values for $\gamma$ include $\{0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$, and for $\tau$ are $\{0.001,0.005,0.01,0.1,1.0,10,100\}$. Their specific choices and impacts across different baseline methods are detailed in Appendix C.
@@
Metrics. We adhere to the suggested evaluation metrics for each benchmark. TOFU adopts two metrics: FQ and MU. FQ measures the extent of data removal by the statistical difference in responses between unlearned models and ground-standard models, which are trained without targeted data. Higher values of FQ are preferred and we report the logarithm of the original FQ values for enhanced readability. MU assesses the overall performance of retention, which is a combination of several more foundational metrics. It can be computed on the retain sets, real authors, and world facts, where higher values indicate better retention.
@@
WMDP performs QA evaluations on WMDP-Bio and WMDP-Cyber to assess the efficacy of removal, where the prompts are standardized following [34]. For retention, WMDP also utilizes QA evaluations but conducted on the MMLU benchmark. Therein, smaller values of QA evaluations are preferred for WMDP-Bio and WMDP-Cyber, while larger values are desired for MMLU. Moreover, MUSE proposes two metrics to assess the removal efficacy, i.e., VerbMem and KnowMem, quantifying various aspects of memorization and membership inference.
MUSE also uses KnowMem for assessing performance retention, where larger values are preferred. To ease analysis, we use the symbols $\uparrow$ and $\downarrow$ next to metric names to indicate that their larger/smaller values are preferred.

### v3

@@ 3.3 Theoretical Analysis
Assume the unlearning objective $\mathcal{L}$ is differentiable, $L$-smooth, and lower bounded. Then, the GRU update rule with the learning rate $\texttt{lr}<{2}/{L}$ will converge to either a) a degenerate configuration where $\cos\bigl{(}\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\bigr{)}=-1$ at a specific step $t$, or b) the locally optimal solution $\bm{\theta}^{*}$ that minimizes $\mathcal{L}(\mathcal{D}_{\mathrm{u}};\bm{\theta})$.
Remark.
Overall, Theorem 3.1 demonstrates that, from a convergence perspective, the GRU does not compromise the original goal of unlearning. This is contingent upon avoiding those cases where $\cos\bigl{(}\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\bigr{)}=-1$. Moreover, given that stochastic optimization is employed for LLM unlearning, we can simply overcome this issue by randomly selecting a new data batch from the unlearning dataset, thereby allowing the unlearning process to continue. Please refer to Appendix A.2 for the detailed proof.
@@
Assume that the retain loss $\mathcal{R}$ is differentiable and $L$-smooth, and the lr-curvature $\mathfrak{H}_{\texttt{lr}}({\mathcal{R}};\bm{g})$ for $\mathcal{R}$ (cf., Definition A.1) satisfies $\mathfrak{H}_{\texttt{lr}}({\mathcal{R}};\bm{g})\geq\ell\|\bm{g}\|^{2}$ for any gradients $\bm{g}$ and some constant $\ell\leq L$. Let $\bm{\theta}^{(t+1)}_{\mathrm{gru}}$ and $\bm{\theta}^{(t+1)}_{\mathrm{u}}$ be the parameters after applying one step of gradient updates for the original $\bm{\theta}^{(t)}$ with and without GRU, respectively. Then, we can ensure $\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta}^{(t+1)}_{\mathrm{gru}})\leq\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta}^{(t+1)}_{\mathrm{u}})$ if a) $\ell\geq L\!\left(1-\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle^{2}/(\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}\,\|\bm{g}_{\mathrm{r}}^{(t)}\|^{2})\right)$ and b) $0<\texttt{lr}\leq\tfrac{2}{L}$.
Remark.
In heuristics, $1-\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle^{2}/(\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}\,\|\bm{g}_{\mathrm{r}}^{(t)}\|^{2}=\sin^{2}\!\phi$ quantifies the degree of conflict between $\bm{g}_{\mathrm{u}}^{(t)}$ and $\bm{g}_{\mathrm{r}}^{(t)}$; larger values (i.e., gradients closer to orthogonal) generally indicate a greater potential to harm the retain performance. Hence, condition a) implies that, when the conflict is more severe, our requirement on the curvature ratio $\ell/L$ must be correspondingly stronger. Condition b) is the classical stability constraint $0<\texttt{lr}\leq 2/L$ for gradient descent on an $L$-smooth function, ensuring the validity of the quadratic bound adopted in GRU. Please refer to Appendix A.3 for the detailed proof.
@@ 4 Go Beyond GRU
Most unlearning methods, including our GRU, rely on retain data to preserve the overall performance. However, the retain data adopted in current benchmarks can often exhibit distributional bias. For example, in the TOFU setup, specific author profiles are selectively unlearned while the remaining profiles are retained. Yet, the broader objective of retention is to preserve model capacity across a diverse range of domains, such as the humanities, sciences, and general knowledge. As a result, the current retain data may not be fully representative, with bias arising from the distributional shift between the adopted retain set and the broader expected data distribution encountered in real-world applications (Huang et al., 2023). It motivates us to investigate a challenging scenario where we need rely exclusively on the unlearn data $\mathcal{D}_{\mathrm{u}}$, without further access to the retain data $\mathcal{D}_{\mathrm{r}}$. To adapt for this setup, we make several adjustments for GRU and propose task vector rectified unlearning (TRU).
@@
The key insight behind TRU is that unlearning typically involves a series of data points rather than a single instance. Thus, for each individual data point $\bm{s}_{\mathrm{u}}\in\mathcal{D}_{\mathrm{u}}$ targeted for unlearning, the remaining data points within $\mathcal{D}_{\mathrm{u}}$, i.e., $\mathcal{D}_{\mathrm{u}}\setminus\{\bm{s}_{\mathrm{u}}\}$, can offer information for retention if used properly. Here, we incorporate the so-called task vectors (Ilharco et al., 2022), which is critical in our algorithmic design.
@@
Task Vector. A task vector typically represents the necessary adjustments for model parameters to incorporate new knowledge. For example, when we want the model to learn from a specific data point $\bm{s}$, we initiate by fine-tuning the current model parameterized, denoted by $\bm{\theta}_{\mathrm{org}}$. It can be achieved through $T$ iterations of gradient updates, following $\bm{\theta}^{(t+1)}=\bm{\theta}^{(t)}+\texttt{lr}\cdot\nabla_{\bm{\theta}}\log p(\bm{s};\bm{\theta}^{(t)})$ with $\bm{\theta}^{(0)}=\bm{\theta}_{\mathrm{org}}$ and $\bm{\theta}_{\bm{s}}=\bm{\theta}^{(T)}$. Obviously, $\bm{T}_{\bm{s}}$ allows for the augmentation of the original model with the knowledge acquired from $\bm{s}$ by applying $\bm{\theta}_{\mathrm{org}}+\bm{T}_{\bm{s}}$. Conversely, to unlearn a data point $\bm{s}_{\mathrm{u}}$, we can reverse this process by subtracting the task vector via $\bm{\theta}_{\mathrm{org}}-\bm{T}_{\bm{s}_{\mathrm{u}}}$ (Barbulescu & Triantafillou, 2024).
@@
Rectified Task Vector. However, the task vector $\bm{T}_{\bm{s}_{\mathrm{u}}}$ still faces the trade-off between unlearning and retention. To address this, we begin by considering a simple scenario to remove a single data point $\bm{s}_{\mathrm{u}}$. Hence, a similar constrained updating rule, as outlined in Eq. (7), can be adopted and further adjusted as:
@@
$&\min_{\tilde{\bm{T}}_{\bm{s}_{\mathrm{u}}}}~{}\|\tilde{\bm{T}}_{\bm{s}_{\mathrm{u}}}-{\bm{T}}_{\bm{s}_{\mathrm{u}}}\|^{2}\\ ~{}\text{s.t.}~{}\langle-\tilde{\bm{T}}_{\bm{s}_{\mathrm{u}}}&,\nabla_{\bm{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\bm{s}_{\mathrm{u}}\};\bm{\theta}_{\mathrm{org}})\rangle\geq 0.\end{aligned}$ (11)
@@
It mandates that the task vector be rectified to have no negative impact on other data points. For this purpose, we utilize the internal reference set for retention, $\mathcal{D}_{\mathrm{u}}\setminus\{\bm{s}_{\mathrm{u}}\}$, to construct a rectified task vector for $\bm{s}_{\mathrm{u}}$. Similar to Eq. (8), we have its closed-form solution as
$\tilde{\bm{T}}_{\bm{s}_{\mathrm{u}}}={\bm{T}}_{\bm{s}_{\mathrm{u}}}$ (12)
@@
$+$ $\frac{[\langle{\bm{T}}_{\bm{s}_{\mathrm{u}}},\,\nabla_{\bm{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\bm{s}_{\mathrm{u}}\};\bm{\theta}_{\mathrm{org}})\rangle]_{-}}{\|\nabla_{\bm{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{u}}\setminus\{\bm{s}_{\mathrm{u}}\};\bm{\theta}_{\mathrm{org}})\|^{2}}\,{\bm{T}}_{\bm{s}_{\mathrm{u}}}.$
@@
This mechanism naturally extends to multiple data points, where we compute a rectified task vector $\tilde{\bm{T}}{\bm{s}\mathrm{u}}$ for each $\bm{s}\mathrm{u}\in\mathcal{D}{\mathrm{u}}$. Moreover, to ensure standardised influence across data points, these task vectors are further normalized so that their magnitudes are equal to 1. The resulting normalized vectors are denoted as $\bar{\bm{T}}_{\bm{s}_{\mathrm{u}}}$, facilitating equitable integration across data points. The final unlearning update is then formed by aggregating (e.g., averaging) these rectified vectors across all elements in $\mathcal{D}_{\mathrm{u}}$:
$\bm{\theta}_{\mathrm{org}}-\frac{\texttt{stg}}{n}\sum_{\bm{s}_{\mathrm{u}}\in\mathcal{D}_{\mathrm{u}}}\bar{\bm{T}}_{\bm{s}_{\mathrm{u}}},$ (13)
@@
where we subtract the average of all normalized task vectors from the original model and $n$ represents the number of data points within $\mathcal{D}_{\mathrm{u}}$ and stg indicates the strength of task vector-based unlearning. It ensures a reliable removal of targeted data while mitigating the compromise to the overall performance. Notably, because each rectified task vector is constructed with respect to its own reference subset, they remain mutually compatible, and their aggregation yields a stable and robust overall unlearning direction. Moreover, when the number of data points $n$ is substantial, an effective strategy involves randomly dividing the entire unlearning set into several smaller batches. Each batch then serves as a substitute for $\bm{s}_{\mathrm{u}}$ in Eq. (11), reducing the demands associated with calculating the task vectors.
@@ 5 Experiments
Benchmarks. Our evaluations adopt three representative benchmarks: TOFU (Maini et al., 2024), WMDP (Li et al., 2024), and MUSE (Shi et al., 2024). TOFU comprises 200 synthetic author profiles, totally 4,000 question-answer pairs. It covers different unlearning setups with varying proportions of data targeted to be unlearned, including 1%, 5%, or 10% of the profiles as unlearning sets. WMDP collects a set of sensitive knowledge encountered in practice, further categorized into three areas as biosecurity, cybersecurity, and chemical security. MUSE constructs their unlearning sets using news articles and books, primarily focusing on addressing copyright issues within existing LLMs.
@@
Baselines and Backbones. For the baseline methods, we focus on a set of objective-based approaches, including GA, GD, NPO, weighted gradient ascent (WGA) (Wang et al., 2024b). All of these methods have demonstrated their practical significance and are thus adopted in our experiments. Moreover, for the backbone models, we adhere to the default suggestions for each benchmarks. We use further fine-tuned LLaMA2-7B-chat (Touvron et al., 2023b) and Phi-1.5 (Abdin et al., 2024) for TOFU; Zephyr-7B-beta (Tunstall et al., 2023) for WMDP; ICLM-7B (Shi et al., 2023) for MUSE.
@@
Hyper-parameters Configurations. In our experiments, We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with the batch size of $32$ and the learning rates $2\times 10^{-5}$ for Phi-1.5 and $1\times 10^{-5}$ for LLaMA2-7B-chat in TOFU; $1\times 10^{-5}$ in MUSE; $4\times 10^{-6}$ in WMDP. Furthermore, the training steps are set to $5$ epochs for TOFU, $1$ epoch for MUSE, and $20$ steps for WMDP. For the hyperparameters within GRU, we employ grid search on validation data to identify their optimal values. The candidate values for $\gamma$ include $\{0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$, and that for $\tau$ are $\{0.001,0.005,0.01,0.1,1.0,10,100\}$. Their specific choices and their impacts across different baseline methods are detailed in Appendix D.
@@
Metrics. We adhere to the suggested evaluation metrics for each benchmark. TOFU adopts two metrics: FQ and MU. FQ measures the extent of data removal by the statistical difference in responses between unlearned models and ground-standard models, which are trained without targeted data. Higher values of FQ are preferred, and we report the logarithm of the original FQ values for enhanced readability. MU assesses the overall performance of retention, which is a combination of several foundational metrics. It can be computed on the retain sets, real authors, and world facts, where higher values indicate better retention.
@@
WMDP performs QA evaluations on WMDP-Bio and WMDP-Cyber to assess the efficacy of removal, where the prompts are standardized following (Gao et al., 2024). For retention, WMDP also utilizes QA evaluations, but conducting on the MMLU benchmark. Therein, smaller values of QA evaluations are preferred for WMDP-Bio and WMDP-Cyber, while larger values are desired for MMLU. Moreover, MUSE proposes two metrics to assess the removal efficacy, i.e., VerbMem and KnowMem, quantifying various aspects of memorization and membership inference. MUSE also uses KnowMem for assessing performance retention, where larger values are preferred. To ease analysis, we use the symbols $\uparrow$ and $\downarrow$ next to metric names to indicate that their larger/smaller values are preferred.